Will somebody help me figure this out and help with formulas
Name: Instructor: Intermediate Accounting. 15 Edition by Kieso, Weygandt, and Warfield Date: Course: P6-4 (Evaluating Payment Alternatives) Howie Long has just learned he has won a 500000 prize in the lottery. The lottery has given him two options for receiving payments: (1) If Howie takes al the money today, the state and the federal governments will deduct taxes at a rate of 46% !immediately (2) Alternatively the lottery offers Howie a payout of 20 |equal payments of Howie will be taxed on each of these payments at a rate of 36,000 with the first payment occurring when Howie turns in the winning ticket 25% Assuming Howie can earn an 8% rate of return (compounded annually) on any money during this period, which pay-out option should he choose? t cash Formula Step 2 Determine the present value of an annuity Percentage Formula Tax burden is IS Text Title
Answer #1

step1: =500000
step 2:
cash payment=500000
tax burden=46%
annual cash yield= cash payment*(1-tax)
=500000*(1-46%)
=270000

other option
monthly payments= amount*(1-tax)
=36000*(1-25%)=27000
use pv formuale in excel
=pv(rate,nper,pmt,fv,type)
=pv(8%,20,27000,0,0)
=265089.98
it is better option 1 than option 2